OpenStudy (anonymous):

What is the correct solution to this problem? A number divided by 7 is 6 less than –8. What is the number? A. –98 B. –56 C. –14 D. –7

OpenStudy (anonymous):

You are given a word problem... now translate the wording to mathematical symbols

OpenStudy (anonymous):

x/7=-8-6 x/7=-14 <-- Multiply both sides by 7 x=-14*7= -98 <--- A